### v1

# Multi-Stage Active Sequential Hypothesis Testing with Clustered Hypotheses

Abstract
We consider the problem where an active Decision-Maker (DM) is tasked to identify the true hypothesis using as few as possible observations while maintaining accuracy. The DM collects observations according to its determined actions and knows the distributions under each hypothesis. We propose a deterministic and adaptive multi-stage hypothesis-elimination algorithm where the DM selects an action, applies it repeatedly, and discards hypotheses in light of its obtained observations. The DM selects actions based on maximal separation expressed by the distance between the parameter vectors of each distribution under each hypothesis. Close distributions can be clustered, simplifying the search and significantly reducing the number of required observations.
Our algorithm achieves vanishing Average Bayes Risk (ABR) as the error probability approaches zero, i.e., the algorithm is asymptotically optimal. Furthermore, we show that the ABR cannot exceed the error probability when the number of hypotheses grows. Simulations are carried out to evaluate the algorithm’s performance compared to another multi-stage hypothesis-elimination algorithm, where an improvement of 5 to 6 orders of magnitude in the mean number of observations required for success is observed.
I Introduction
HT is a statistical method used to evaluate competing hypotheses and determine which one best fits a sampled population based on sample data. HT has been in use for centuries, long before formal statistical methods were developed, and has been applied across highly diverse domains. For example, it has been used to identify the physical model that best fits a set of measurements, such as determining the Earth’s circumference or verifying a quantum mechanics hypothesis. HT is also an integral part of modern scientific practice, particularly in medical research where it’s used to evaluate the effectiveness of treatments and diagnose conditions based on clinical tests (e.g., testing for infectious diseases like COVID-19).
In recent years, HT has gained additional significance with the recent vast interest in Machine Learning (ML) and Artificial Intelligence (AI). Modern computing systems, including autonomous control systems, anomaly detection algorithms, quality control systems, and networked sensors, rely on various HT techniques to detect events of interest or classify specific occurrences. This detection or classification task is typically performed using a Decision-Maker (DM), which determines which hypothesis (from two or more alternatives) is best supported by the available data. In classic binary HT, this is often accomplished by computing the LLR or its non-logarithmic equivalent, comparing it to a predetermined threshold, and selecting the corresponding hypothesis based on this comparison.
In many practical scenarios, the DM needs to collect and process data in real-time, and often can select which samples or data sources to examine next. For example, a physician can determine which subsequent diagnostic tests a patient should undergo to identify their condition, or a network administrator can choose which routers to monitor to detect and classify potential cyber-attacks. This type of sequential decision-making about data collection allows for more efficient and targeted information gathering. In such settings, the DM must balance two central objectives: maximizing decision accuracy while minimizing the time delay (typically measured by the number of observations) until a decision is reached.
In his seminal work [1], Wald addressed the challenge of sequential data acquisition in a decision-making setup involving two competing hypotheses by introducing binary Sequential (SHT). The proposed scheme, which we term the Wald Test in the sequel, establishes two thresholds, each corresponding to one of the hypotheses. The DM sequentially collects observations and updates the LLR until it exceeds one of the thresholds, at which point the corresponding hypothesis is declared correct. Wald’s results demonstrated that SHT allows the DM to reach decisions more quickly than traditional fixed-sample hypothesis testing while maintaining the same level of accuracy.
The Wald test has been generalized and expanded over the years. Notably, the Armitage test [2, 3] extends the Wald test to sequential multi-hypothesis testing by conducting “tournaments” between hypotheses, where each drawn sample serves as part of a Wald test “contest” between any two hypotheses. In this context, the DM declares the tournament winner as the underlying hypothesis.
Chernoff [4] extended the Wald test setup to incorporate actions, leading to active binary SHT. In Chernoff’s model, the DM can choose from a set of actions, each yielding an observation that follows a specific known distribution. For example, to diagnose a patient’s illness, a doctor can select which examination the patient should undergo next, where the distribution of results for each possible illness depends on the chosen examination. Chernoff proposed a stochastic policy (i.e., actions are drawn according to some non-degenerate distribution) for selecting actions sequentially, allowing the DM to “shape” observations through chosen actions to optimize accuracy and detection delay. Chernoff’s stochastic approach was later fine-tuned by Bessler in [5] and by Nitinawarat et al. in [6] to encompass active multihypothesis testing like considered in this work. Solutions based on Dynamic Programming and Linear Fractional Programming also exist, e.g., [7, 8, 9]. Deep Learning and Reinforcement Learning have also been considered, e.g., [10, 11, 12].
A prime example of a deterministic action-selecting policy is the DGF policy used for anomaly detection in [13] and its numerous extensions, e.g., [14, 15, 16]. In the DGF policy, one observes that the accumulated LLRs corresponding to infected and clean distributions are random walks growing apart when enough observations are collected and when the gap between the largest and second-largest LLRs is sufficiently large, the DGF policy stops and declares the hypothesis corresponding to the highest accumulated LLR as true. Although the DGF policy outperforms Chernoff’s approach, the main drawback of the DGF policy is that it is tailored for anomaly detection. Namely, it assumes actions produce observations from only two possible distributions.
In this paper, we take a different approach from the previously mentioned works by adopting the elimination approach rather than the typical search approach. Namely, instead of looking for enough evidence to conclude which hypothesis is true, the obtained observations are utilized to identify which hypothesis is “definitely not true.” To this end, the DM leverages the ability of some actions to separate some hypotheses better than others to swiftly and accurately discard inconsistent hypotheses. In particular, the DM selects its current action based on maximal separation of the remaining hypotheses, applies it repeatedly to collect enough observations to infer which hypotheses are incorrect using the Armitage Test, and discards them before repeating the procedure unless a single hypothesis remains.
Still, it is possible for the maximal separation approach to require the DM to collect a large number of observations, all the more so when the obtained observations follow highly similar distributions under different hypotheses. Thus, we suggest clustering these distributions and require the DM only to identify which cluster is correct and discard all incorrect hypotheses accordingly. The closest work to our approach is the recently proposed algorithm by Gan, Jia, and Li in [17, Algorithm 2], which we refer to as the GJL algorithm. In GJL, the action capable of minimizing the number of surviving hypotheses is selected in each stage, and the (fixed) number of observations is based on separating the two closest, yet not identical, observation distributions. This philosophy of minimal survivors directly contrasts our approach of fast elimination incited by the maximal separation policy. Assuming that the clustering is done properly, we show that our algorithm is asymptotically optimal in terms of vanishing Average Bayes Risk (ABR) when the desired error probability approaches zero and is bounded when the number of hypotheses grows. Finally, we numerically evaluate our algorithm and observe a significant performance improvement compared to GJL. Specifically, our algorithm improves the mean number of observations required until termination (and consequently, the ABR) by 5 to 6 orders of magnitude for any desired error probability.
II System Model
II-A Notation
All vectors in this manuscript are underlined (e.g., ${\underline{x}}$). The $\ell_{2}$-norm between two vectors ${\underline{x}}$ and ${\underline{y}}$ is denoted as $\|{\underline{x}}-{\underline{y}}\|_{2}$. For the Kullback-Leibler Divergence (KLD) between two distributions, $f$ and $g$, we write $\mathcal{D}_{KL}(f\|g)$. Unless explicitly specified (e.g., $\ln$), all logarithms in this manuscript are in base two. Throughout this paper, we adopt the Bachmann–Landau big-O asymptotic notation, $\operatorname{\mathcal{O}}\left(\cdot\right)$, as defined in [18, Chapter 3].
II-B Model
The system model consists of a single, active, DM capable of obtaining observations according to the actions taken. The DM is tasked with correctly identifying the underlying hypothesis out of a finite set of hypotheses =$\mathcal{H}=\{0,1,\dots H-1\}$. The prior probability of hypothesis $i$ (or $H_{i}$ for short) is $\pi_{i}$, where $10<\pi_{i}<$ 1 for all $i$ to avoid triviality.
Let $\mathcal{A}\subset\mathbb{N}$ with $|\mathcal{A}|<\infty$ be the set of all actions available to the DM. We assume that the observations are independent and follow the same distribution when action $a_{n}$ is taken, with the distribution parameters depending on the underlying hypothesis. Namely, assuming $H_{i}$ is the underlying hypothesis and that action $a_{n}\in\mathcal{A}$ is taken at time step $n$, the DM obtains $X_{n}\sim f_{a_{n}}(\cdot\ ;{\underline{\theta}}_{i}(a_{n}))$, where ${\underline{\theta}}_{i}(a_{n})\in\mathbb{R}^{M_{a_{n}}}$ is the (vectorized) distribution parameter under $H_{i}$ and $f_{a_{n}}$ is its Probability Density Function (PDF). Extension to non-scalar samples is straightforward and will not be discussed in this paper. All distributions and their parameters under each and every hypothesis are assumed to be known by the DM. Figure 1 visualizes the model.
Figure 1: System model. The DM is tasked to identify the correct hypothesis (say $H_{i}$) out of $H$ possible hypotheses. By taking action $a_{n}$ at time step $n$, the DM obtains an observation $x_{n}$ whose PDF is $f_{a_{n}}(\cdot\ ;{\underline{\theta}}_{i}(a_{n}))$. The alphabet and size of $x_{n}$ and ${\underline{\theta}}_{i}(a_{n})$ may depend on the action $a_{n}$.
Like [17], we make three additional assumptions:
  1. (A1)
    (Seperation) For any $a\in\mathcal{A}$ and $i$, $j\in\mathcal{H}$, $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}$ is either 0 or greater than some $\alpha>0$. Furthermore, there is no $a$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}=0$ for all $i$, $j\in\mathcal{H}$. Note that $\alpha$ can be very small.
  2. (A2)
    For any action $a$ there are $0<c_{1}\leq c_{2}$ such that any $i,j\in\mathcal{H}$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}>0$ also have $c_{1}\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{j}(a))\|f_{a}(\cdot;{\underline{\theta}}_{i}(a)))\leq\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}\leq c_{2}\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{j}(a))\|f_{a}(\cdot;{\underline{\theta}}_{i}(a)))\leq c_{2}\beta$ for some $\beta$.
  3. (A3)
    (Validity) For all $i,j\in\mathcal{H}$ with $i\neq j$, there is some $a\in\mathcal{A}$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}>0$.
The first assumption ensures that there are no meaningless actions and that distribution parameters are separated in $\mathbb{R}^{M_{a}}$. The second is technical and allows us to simplify hypothesis comparison by using the $\ell_{2}$-norm on the parameters rather than computationally prohibitive metrics on distributions (e.g., Total Variation Distance [19, Chapter 13.1]). The third assumption assures that multi-stage hypothesis-eliminating algorithms can stop.
Let $\Phi$ be the source selection process generating the action sequence $\{A_{n}\}_{n=1}^{\infty}$. The source selection rule is non-adaptive if the actions do not depend on the gathered data for any time step and is adaptive otherwise. It may also be either deterministic or stochastic. The decision rule is given by $\Psi\in\mathcal{H}$, i.e., $\Psi=i$ implies that $H_{i}$ is declared as true.
Like many other works in the literature, e.g., [4, 6, 13, 20, 16], we assume the DM pays $\delta\in(0,1)$ to acquire a realization of $X_{n}$, and we are interested in the asymptotic behavior when $\delta\to 0$. Additionally, to emphasize the gains from the hypothesis clustering (which will be discussed in the following section), we are also interested in the regimes when $H$ is not too small, and, accordingly, we modify the common Bayesian approach. Namely, instead of the unit penalty accepted in other works, we assign a penalty of $H^{2}$ for a wrong declaration.
Let $\Gamma\triangleq(\Phi,\Psi)$ be an admissible strategy for the SHT aiming at minimizing the ABR, $\delta\mathbb{E}\left[N\right]+H^{2}\times p_{e}$, where $N$ is the number of observations upon algorithm termination (i.e., detection delay), and $p_{e}$ is the error probability of the strategy $\Gamma$. Equivalently, it is a strategy minimizing
$\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}.$ (1)
Notably, (1) weights two key components of the problem: the error probability, $p_{e}$, and the mean detection delay $\mathbb{E}\left[N\right]$. By increasing the number of samples, the DM can significantly reduce its error probability, but increasing it too much makes the first addend the dominant penalty. Conversely, taking too few samples will increase the error probability, making it the dominant penalty. Accordingly, the desired error probability and the mean detection delay must be balanced. The normalized cost, $\frac{\delta}{H^{2}}$, plays a role in damping the mean detection delay; Specifically, it vanishes in $\delta$ only if the mean detection delay is sublinear in $\frac{1}{\delta}$ and subquadratic in $H$. While the motivation behind the former choice is clear, the latter choice damps the number of samples per stage to be sublinear in $H$ in cases when $H-1$ stages are used in a multi-stage hypothesis-eliminating algorithm.
III Observation Acquisition Policy
In this section, we formalize the suggested algorithm to acquire observations and compute the underlying hypothesis using them while minimizing the ABR. Like many works in literature, in order to identify which hypothesis is correct or not, the DM leverages the Test (LLRT) where the collected samples up to time step $n$ are used to compute the accumulated LLR which is compared to some fixed threshold $\gamma>0$ we will discuss later. Namely, if the DM uses only action $a$, the pair-wise per-action accumulated LLR is:
$L_{ij}(a,n)$ $\triangleq\sum_{t=1}^{n}\log\frac{f_{a}(x_{t};{\underline{\theta}}_{i}(a))}{f_{a}(x_{t};{\underline{\theta}}_{j}(a))}.$
For simplicity, we will drop $a$ and $n$ and write $L_{ij}$. In our context, $H_{i}$ wins against $H_{j}$ if $L_{ij}\geq\gamma$, loses against $H_{j}$ if $L_{ij}<-\gamma$, and competes against $H_{j}$ otherwise. We will show in the following section that $2^{-\gamma}$ controls the per-stage error probability of the algorithm.
Generally, identifying the underlying hypothesis has two distinct approaches. The first is the classic and straightforward approach common in literature where the DM collects observations and computes $L_{ij}$ until some $i^{*}$ wins against all other hypotheses (e.g., [2, 5]). While this approach is simple and fast, it may fail when the environment outputs observations with identical distributions under many different hypotheses, or in other words, the case we are interested in solving.
The second approach, tailored for the above case, is hypothesis eliminating where the DM identifies which hypotheses are “definitely not true” and discards them until a single hypothesis remains and is declared as the underlying hypothesis. Notably, algorithms of the first approach can be converted to the second by discarding the losing hypotheses and retrying using a different set of actions, as we do in this work.
Once an action is selected, eliminating some hypothesis is straightforward by finding at least one hypothesis that wins against it. Thus, a multi-stage algorithm based on the LLRT is sufficient for one-by-one hypothesis elimination in our case since there is always some action capable of separating at least one hypothesis from the others due to assumption (A3). Naturally, one-by-one elimination is inefficient but can be improved by the action selection policy. Namely, at each stage, the DM can greedily pick the action capable of maximizing the number of losing hypotheses like in GJL [17, Algorithm 2]. However, such a policy may result in a large number of samples and, accordingly, in large ABR.
We adopt a different policy with both efficient hypothesis elimination (i.e., eliminating a set of hypotheses in each stage) and vanishing ABR. In each stage, we compute the action maximizing the separation, thus minimizing the number of samples required to eliminate hypotheses and, finally, minimizing the ABR. To enhance the hypothesis elimination process, we further suggest clustering hypotheses (per action) by proximity, as we will elaborate soon. Here, the DM only needs to identify the correct cluster rather than a set of hypotheses whose output distribution is identical and, accordingly, may use fewer samples.
III-A Preprocessing Step - Per-Action Hypothesis Clustering
In this subsection, we explain the hypothesis clustering concept in depth. The main purpose of the hypothesis clustering is to allow the DM to refine its “bad actions,” characterized by small KLD between some of their output distributions (e.g., either zero or very close to $\alpha$ from assumption (A1)). Namely, to tell apart between two or more very close hypotheses, the DM would require, on average, a large number of observations. Therefore, since these hypotheses are very similar, the DM can significantly reduce the number of the collected observations by simply identifying the set of similar hypotheses (e.g., by identifying a single member of these hypotheses as we do in this work) and using a different action with better separation on the found set in the next stage. In other words, the DM can improve its performance in terms of mean detection delay (and, consequently, obtain lower ABR) by clustering hypotheses by their informational proximity defined by a proximity parameter $\varepsilon$ set by the user.
Hypothesis clustering by their proximity (i.e., selecting $\varepsilon$) and selecting a representative from each cluster are not trivial. Thus, we first discuss a key property of a good hypothesis clustering mechanism. Assume action $a$ is repeatedly used and that $H_{i}$ is the underlying hypothesis. Now, we observe the normalized accumulated LLR between $H_{k}$ in $H_{i}$’s cluster against $H_{j}$ that is not in $H_{i}$’s cluster converges to
$\frac{1}{n}L_{kj}$ $\to\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{k}(a))}{f_{a}(X;{\underline{\theta}}_{j}(a))}\middle|H_{i}\right]$
$=\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{j}(a)))$
$\quad\quad-\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{k}(a)))\triangleq\Delta\mathcal{D}_{ijk}$ (2)
in probability from the Asymptotic Equipartition Property (AEP) property [21, Theorem 11.8.1]. Since $\frac{1}{n}\gamma\to 0$ for any fixed $\gamma$, we observe that $H_{i}$ can be declared as true if and only if the difference in (2) is positive. Therefore, any $H_{k}$ clustered with $H_{i}$ should have its KLD (from $H_{i}$ to $H_{k}$) strictly lower than $H_{i}$’s KLD from any $H_{j}$ not clustered with $H_{i}$.
Since KLD is asymmetric, $H_{i}$ can be in the same cluster as $H_{k}$, but not vice-versa. To circumvent this issue, we leverage assumption (A2) that bounds the KLD using a symmetric metric property, the $\ell_{2}$-norm between the output distribution parameters for appropriate $c_{1}$ and $c_{2}$. Specifically, the use of the $\ell_{2}$-norm allows us to rewrite the property in (2) as $c\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}>\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{k}(a)\|_{2}^{2}$ for any $H_{k}$ in $H_{i}$’s cluster and for any $H_{j}$ not in $H_{i}$’s cluster for some $c$.
Now, we can formalize the desired clustering on the hypotheses using their parameter space. Let $\operatorname{param}\left(a\right)\triangleq\{{\underline{\theta}}_{i}(a):i\in\mathcal{H}\}$ be the set of distribution parameters under action $a$. Let $\mathcal{B}_{\varepsilon}({\underline{\nu}})\triangleq\{{\underline{\theta}}\in\operatorname{param}\left(a\right):\|{\underline{\theta}}-{\underline{\nu}}\|_{2}^{2}\leq\varepsilon\}$ be the $\varepsilon$-neighborhood of ${\underline{\nu}}$ composed of points from $\operatorname{param}\left(a\right)$. We say that ${\underline{y}}$ is reachable from ${\underline{x}}$ if there are ${\underline{x}}_{1},{\underline{x}}_{2},\dots,{\underline{x}}_{r-1}$ such that ${\underline{x}}={\underline{x}}_{1}$, ${\underline{x}}_{j}\in\mathcal{B}_{\varepsilon}({\underline{x}}_{j+1})$ and, additionally we have ${\underline{x}}_{r-1}\in\mathcal{B}_{\varepsilon}({\underline{y}})$. Similarly, ${\underline{x}}$ and ${\underline{y}}$ are connected if ${\underline{x}}$ is reachable from ${\underline{y}}$ and vice-versa.
Thus, we define the cluster of $H_{i}$ as $\mathcal{C}_{i}(a,\varepsilon)\subseteq\operatorname{param}\left(a\right)$ to be a non-empty subset such that (i) for any ${\underline{\nu}}$, if ${\underline{\nu}}$ is reachable from ${\underline{\theta}}_{i}$ then ${\underline{\nu}}\in\mathcal{C}_{i}(a,\varepsilon)$, and (ii) any ${\underline{\nu}},{\underline{\theta}}\in\mathcal{C}_{i}(a,\varepsilon)$ are connected. Now, we define
$\operatorname{equiv}\left(i,a\right)$$\triangleq\left\{k\in\mathcal{H}\ :\ {\underline{\theta}}_{k}\in\mathcal{C}_{i}(a,\varepsilon)\right\}$
$\operatorname{repr}\left(\mathcal{U},a\right)$$\triangleq\left\{\min\{\mathcal{U}\cap\operatorname{equiv}\left(i,a\right)\}:i\in\mathcal{U}\right\}$
to be the set of equivalent hypotheses to $H_{i}$ and the set of representatives taken from $\mathcal{U}\subseteq\mathcal{H}$, respectively. Note that the choice of the smallest indices is arbitrary. If no clustering is used, we define $\operatorname{equiv}\left(i,a\right)\triangleq\{j\in\mathcal{H}:{\underline{\theta}}_{j}(a)={\underline{\theta}}_{i}(a)\}$ instead. Remark: We assume that even after the clustering, assumption (A3) holds.
The use of the $\ell_{2}$-norm allows us to use rich density-based clustering over $\mathbb{R}^{M_{a}}$ (e.g., [22, Chapter 5]) to cluster hypotheses by proximity. Notably, the famous Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [23] is tailored to our needs and is sufficient for clustering. DBSCAN groups together points with many shared neighbors (specified by the MinPts parameter) within their $\varepsilon$-neighborhood and labels all remaining points as noise or outliers. Furthermore, DBSCAN guarantees that clusters are $\varepsilon$ apart from each other.
III-B Multi-Staged Algorithm
In this subsection, we present and describe our multi-stage algorithm. The procedure is given in Algorithm 1. In Line 1, we initialize the “alive” hypotheses, $H_{\mathrm{alive}}$, to be the set of all hypotheses. The algorithm thins this set in each stage until one or no hypothesis remains. If a single hypothesis remains at the end of the procedure, it is declared as true, and an error is raised otherwise.
We term Lines 3-11 as stages. In each stage, an action $a$ is computed (Line 3). Specifically, each action is associated with a non-zero minimal distance between each two alive hypotheses, and the action with the maximal minimum distance is selected for the current round. Line 4 computes the representatives of each cluster out of the alive hypotheses. Now, we apply action $a$ repeatedly until a winner against all hypotheses, $i^{*}$, is found (Lines 6-9). Then, we update $H_{\mathrm{alive}}$ to be the alive hypotheses in $i^{*}$’s cluster (Lines 10-11).
Algorithm 1 Multi-Stage LLR-Based SHT
$H_{\mathrm{alive}}\leftarrow[0,H-1]$triangleright Initialize alive hypotheses
2:while  $|H_{\mathrm{alive}}|\geq 2$  do
3:     $a\leftarrow\displaystyle\mathop{\operatorname{argmax}}_{b\in\mathcal{A}}\left\{\min_{\begin{subarray}{c}i,j\in H_{\mathrm{alive}}\\ j\not\in\operatorname{equiv}\left(i,b\right)\end{subarray}}\|{\underline{\theta}}_{i}(b)-{\underline{\theta}}_{j}(b)\|_{2}^{2}\right\}$
4:     $\tilde{H}\leftarrow$ $\operatorname{repr}\left(H_{\mathrm{alive}},a\right)$
5:     $L_{ij}\leftarrow 0\ \forall i\neq j\in\tilde{H}$triangleright Initialize LLRs
6:     while  $\not\exists i\in\tilde{H}$: $L_{ij}\geq\log\frac{H}{\delta}\ \forall j\neq i$  do
7:         acquire $x$ according to action $a$
8:         $L_{ij}\leftarrow L_{ij}+\log\frac{f_{a}(x;{\underline{\theta}}_{i}(a))}{f_{a}(x;{\underline{\theta}}_{j}(a))}$ for all $i\neq j\in\tilde{H}$
9:     end while
10:     $\hat{H}\leftarrow\operatorname{equiv}\left(i^{*},a\right)$triangleright $i^{*}$ has $L_{i^{*}j}\geq\log\frac{H}{\delta}\ \forall j$
11:     $H_{\mathrm{alive}}\leftarrow H_{\mathrm{alive}}\cap\hat{H}$triangleright Update alive hypotheses
12:end while
13:return $H_{\mathrm{alive}}$
IV Analysis
In this section, we analyze Algorithm 1 in terms of error probability, mean detection delay, and the scaling of the ABR as functions of $\delta$ and $H$. Notably, all previous results regarding the Armitage Test, e.g., vanishing error probability or bounded mean detection delay, do not trivially follow from using the Armitage Test at each stage due to the clustering, and accordingly, all computations must be modified. Due to brevity, the proofs are attached in Appendix A.
We start with the error probability. To bound the error probability of the algorithm, we first bound the per-stage error probability. Each stage of the algorithm is terminated when some hypothesis has all of its LLRs exceed $\log\frac{H}{\delta}$. At first glance, there are two possible error events; the first is that the stage winner (at some stage) is not the representative of the real hypothesis, and the second is the event when two different hypotheses win in the current round. We argue that the latter case never occurs in the following lemma:
Lemma 1.
Let $i_{r}^{*}$ be the winner at stage $r$. Then, $i_{r}^{*}$ is unique.
Proof:
See Appendix A-A. ∎
Now that we have established that there are no ties, we bound the former error probability with the following lemma:
Lemma 2.
For any action $a$, there exists some $\eta_{a}\in[0,1)$ such that $\Delta\mathcal{D}_{ijk}>\eta_{a}$, the current action still can distinguish between at least two hypotheses, and for any $\varepsilon\leq\eta_{a}$, the per-stage error probability cannot exceed $\frac{\delta}{H}$. Particularly, $p_{e}\leq\delta$.
Proof:
See Appendix A-B. ∎
Now, we focus on the scaling laws of the mean detection delay, $\mathbb{E}\left[N\right]$, given in the following lemma:
Lemma 3.
Let $\tau_{r}$ be the number of observations used in stage $r$ when using action $a_{r}$, $H_{\mathrm{alive}}^{(r)}$ be the alive hypotheses in stage $r$, and $k$ be the representative from $H_{i}$’s cluster. Then $\mathbb{E}\left[\tau_{r}|H_{i}\right]\approx\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ for any $i$. Furthermore, $\mathbb{E}\left[\tau_{r}|H_{i}\right]$ is bounded, i.e., $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}|H_{i}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$ and particularly, $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}$.
Proof:
See Appendix A-C. ∎
Remark: Although $\mathbb{E}\left[N\right]$ seems identical to non-clustering cases, it is, in fact, smaller as the expectation not only contains fewer positive addends, but each addend is also smaller.
Corollary 1 (Asymptotic Optimality).
For a suitable choice of $\varepsilon$, Algorithm 1 has vanishing ABR when $\delta\to 0$. Furthermore, the ABR does not exceed $\delta$ when $H\to\infty$.
Proof:
See Appendix A-D. ∎
V Numerical Results
In this section, we present numerical results illustrating the performance of our algorithm compared to GJL. To this end, we have conducted simulations. Namely, we compare GJL to our algorithm in two scenarios: (i) when no clustering is used. (ii) clustering is used with a suitable proximity parameter.
In the simulations, we have used uniform priors, i.e., $\pi_{i}=\frac{1}{H}$. For a fair comparison, we picked distributions whose mean can distinguish between different hypotheses, namely, the exponential and unit-variance normal distributions. The number of hypotheses was set to $H=16$, and the DM had $|\mathcal{A}|=17$ actions. The first $H$ actions had a mean of 3 if $H_{a-1}$ is true for $1\leq a\leq H$, and a random mean drawn uniformly from $[0,1]$ otherwise. The last action had the means follow the equation $0.5+0.01\times i$ for $0\leq i\leq H$. All random means were drawn only once. For the clustering algorithm, DBSCAN, we use $\text{minPts}=1$ and proximity parameters $\varepsilon\in\{0,0.1\}$.
In Figure 2, we present the results for the two cases. The dashed lines present the ABR for the case where all actions produce observations following the normal distribution, whereas non-dashed lines present the ABR for the case where all actions produce exponentially distributed observations. The orange curves correspond to GJL, and the blue and green curves correspond to our algorithm without and with clustering, respectively. Each algorithm attains vanishing ABR as $\delta\to 0$, but the non-greedy selections made by each instance of our algorithm enjoys significantly lower ABR, which is attributed to lower mean detection delay.
[…]
Figure 2: Comparison of the ABR (equation (1)) for $H=16$ using our algorithm (with proximity parameter $\varepsilon\in\{0,0.1\}$) against GJL. The observations obtained are either normally distributed (dashed) or exponentially distributed (non-dashed). The ABR vanishes as $\delta\to 0$ for all instances, but GJL’s greedy selection has a slower decay.
Appendix A Miscellaneous Proofs
A-A Proof of Lemma 1
Assume to the contradiction that there are two winners, $i_{r}^{*}$ and $j_{r}^{*}$. Since $i_{r}^{*}$ is a winner, then $L_{i_{r}^{*}j_{r}^{*}}\geq\log\frac{H}{\delta}$. Likewise, $j_{r}^{*}$ is a winner so $L_{j_{r}^{*}i_{r}^{*}}\geq\log\frac{H}{\delta}$. However, due to symmetry, we have $\log\frac{H}{\delta}\leq L_{i_{r}^{*}j_{r}^{*}}=-L_{j_{r}^{*}i_{r}^{*}}\leq-\log\frac{H}{\delta}$, contradiction.
A-B Proof of Lemma 2
Since the number of alive hypotheses can only decrease in each stage, there can be no more than $H$ stages. Consequently, we can compute the error probability of a single stage, and then, by the union bound, the error probability of the algorithm, $p_{e}$, is at most $H$ times the error probability of a single stage. Hence, we will focus on computing the first stage’s error probability for simplicity.
Here, we write ${\underline{X}}_{1}^{M}$ as a shorthand notation for $(X_{1},X_{2},\dots,X_{M})^{T}$, and, with slight abuse of notation, $f_{a}({\underline{x}}_{1}^{M};{\underline{\theta}}_{l}(a))$ is their PDF (product distribution) for any $l$ under action $a$. Let $\operatorname{\mathbb{P}}\left(j|i\right)$ be the probability that $H_{j}$ erroneously survives a stage when $H_{i}$ is the underlying hypothesis, and, consequently, $H_{i}$ does not survive the current stage. Let $a$ be the action the DM takes in the current stage. Let $k$ be the representative from $H_{i}$’s cluster. Let $\mathds{1}{\left\{A\right\}}$ be the indicator for the event $A$. Let $\mathcal{R}_{j}$ be the decision region of hypothesis $H_{j}$, i.e. if $\tau$ is the stopping time of the stage (due to exceeded LLR), then the DM declares $H_{j}$ as true if the observation sequence ${\underline{x}}_{1}^{\tau}\in\mathcal{R}_{j}$. Therefore, by definition,
$\operatorname{\mathbb{P}}\left(j|i\right)$ $=\mathbb{E}\left[\mathds{1}{\left\{{\underline{X}}_{1}^{\tau}\in\mathcal{R}_{j}\right\}}|H_{i}\right].$ (3)
Upon round termination, since $H_{j}$ survives then for any $l\in\operatorname{repr}\left(H_{\mathrm{alive}},a\right)$, and particularly for $k$, we have
$\sum_{t=1}^{\tau}\log\frac{f_{a}(x_{t};{\underline{\theta}}_{j}(a))}{f_{a}(x_{t};{\underline{\theta}}_{l}(a))}\geq\log\frac{H}{\delta}$
for any ${\underline{x}}_{1}^{\tau}\in\mathcal{R}_{j}$. Thus, for any ${\underline{x}}_{1}^{\tau}\in\mathcal{R}_{j}$ we have
$\prod_{t=1}^{\tau}f_{a}(x_{t};{\underline{\theta}}_{k}(a))\leq\frac{\delta}{H}\prod_{t=1}^{\tau}f_{a}(x_{t};{\underline{\theta}}_{j}(a)).$ (4)
Now we bound (3) as follows:
$\operatorname{\mathbb{P}}\left(j|i\right)$ $=\mathbb{E}\left[\mathds{1}{\left\{{\underline{X}}_{1}^{\tau}\in\mathcal{R}_{j}\right\}}|H_{i}\right]$
$=\mathbb{E}\left[\mathds{1}{\left\{{\underline{X}}_{1}^{\tau}\in\mathcal{R}_{j}\right\}}\frac{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{k}(a))}{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]$
$\leq\frac{\delta}{H}\mathbb{E}\left[\mathds{1}{\left\{{\underline{X}}_{1}^{\tau}\in\mathcal{R}_{j}\right\}}\frac{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{j}(a))}{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]$
$\leq\frac{\delta}{H}\mathbb{E}\left[\frac{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{j}(a))}{f_{a}({\underline{X}}_{1}^{\tau};{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]$
$=\frac{\delta}{H}\left(\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{j}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]\right)^{\tau}.$
The first inequality follows (4), and the second discards the indicator. The last transition follows the i.i.d. assumption on the obtained observations. We are left to show that the expectation powered by $\tau$ is less than or equal to 1.
Now, instead of taking the expectation with respect to $f_{a}(X_{1};{\underline{\theta}}_{i}(a))$, we take it with respect to $f_{a}(X_{1};{\underline{\theta}}_{j}(a))$ to obtain:
$\operatorname{\mathbb{P}}\left(j|i\right)\leq\frac{\delta}{H}\left(\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]\right)^{\tau}.$
Although the change of measure did not change the expectation powered by $\tau$, it is easier (conceptually) to bound the likelihood ratio between two PDFs of hypotheses in the same cluster rather than the likelihood ratio between two PDFs of hypotheses from different clusters. Observe that
$-\Delta\mathcal{D}_{ijk}$ $=\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{j}(a))}{f_{a}(X;{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]$
$\leq\frac{1}{\ln 2}\left(\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{j}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{i}\right]-1\right)$
$=\frac{1}{\ln 2}\left(\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]-1\right)$
from the famous inequality $\ln x\leq x-1$. Thus,
$\Delta\mathcal{D}_{ijk}$ $\geq\frac{1}{\ln 2}\left(1-\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]\right).$
In order to have $\Delta\mathcal{D}_{ijk}>\eta_{ai}$ for some $\eta_{ai}\in[0,1)$, it is sufficient to show that there exists some $\eta_{ai}$ such that
$1-\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]\geq\eta_{ai}$
for any $H_{k}$ in $H_{i}$’s cluster and $H_{j}$ not in $H_{i}$’s cluster. However, such $\eta_{ai}$ always exists since $H_{i}$ is allowed to be isolated and represented by itself, forcing $\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]=1$. Otherwise, if $H_{i}$ is not isolated and we have
$\mathbb{E}\left[\frac{f_{a}(X_{1};{\underline{\theta}}_{i}(a))}{f_{a}(X_{1};{\underline{\theta}}_{k}(a))}\middle|H_{j}\right]\leq 1-\eta_{ai}\leq 1.$
Thus, $\eta_{a}=\min_{i}\eta_{ai}$ is sufficient and $\operatorname{\mathbb{P}}\left(j|i\right)\leq\frac{\delta}{H}$ as desired. Now, instead of using the conservative computation (select $\operatorname{\mathbb{P}}\left(\text{select }j\neq i|i\right)=\sum_{j\neq i}\operatorname{\mathbb{P}}\left(j|i\right)$ in order to bound the per-stage error probability (cf. [2, Equation (4)]), we instead use the following observation: $H_{j}$ is the stage winner only after winning against all other hypotheses in all $\operatorname{\mathcal{O}}\left(H^{2}\right)$ Wald Tests, so the error probability in the current stage is the probability that $H_{j}$ is accepted in the Wald Test composed of comparing $H_{j}$ and $H_{i}$. Consequently, $\frac{\delta}{H}$ bounds the per-stage error probability, and as the section preamble suggests, the error probability of the algorithm cannot exceed $\frac{\delta}{H}\times H=\delta$.
A-C Proof of Lemma 3
Much like in the proof of the previous lemma, we notice that the Armitage Test essentially conducts $\operatorname{\mathcal{O}}\left(H^{2}\right)$ Wald Tests in parallel using the same observation vector. Due to the algorithm’s stage termination condition when clustering, we are interested in $\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ and the number of observations required for it to exceed $\log\frac{H}{\delta}$, where $k$ is the index of the hypothesis representing $H_{i}$ like before. Using Wald’s Identity, we have
$\mathbb{E}\left[L_{kj}|H_{i}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{\tau_{r}}\log\frac{f_{a_{r}}(x_{t};{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(x_{t};{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\mathbb{E}\left[\log\frac{f_{a_{r}}(X;{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(X;{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\Delta\mathcal{D}_{ijk}$ (5)
for any relevant $i$, $k$, and $j$ (and particularly, for the LLR minimizer). Denote $\hat{p}_{e}$ as the per-stage error probability. From the stopping rule, we have
$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]=\hat{p}_{e}\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\exists j\neq i\text{ wins}\right]$
$+(1-\hat{p}_{e})\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\min_{l}L_{kl}\geq\log\frac{H}{\delta}\right]$
using Bayes. Now, in similar arguments made by Wald in [1, Chapter 4], we approximate $\hat{p}_{e}\approx\frac{\delta}{H}$ for small enough $\delta$. Furthermore, conditioned on success, we approximate $\min_{l}L_{kl}\approx\log\frac{H}{\delta}$ for small enough $\delta$. On the other hand, conditioned on error, we observe that $\min_{l}L_{kl}<\log\frac{H}{\delta}$ since otherwise there would be no error as the winner is unique per Lemma 1. This inequality can be weakened, and we obtain that
$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ $\lesssim\left(1-\frac{\delta}{H}\right)\log\frac{H}{\delta}+\frac{\delta}{H}\log\frac{H}{\delta}$
$=\log\frac{H}{\delta}.$ (6)
Specifically, combining (5) and (6) yields
$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\lesssim\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ (7)
and the upper bound follows. Finally, since $i$ and $j$ are not clustered together, $\Delta\mathcal{D}_{ijk}\geq\varepsilon$ and we bound (7) by $\frac{1}{\varepsilon}\log\frac{H}{\delta}$. The lower bound
$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\geq\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$
$\geq\frac{\log\frac{H}{\delta}}{\beta}$
follows from applying [24, Theorem 3.1] (using the law of large numbers as in (2)) as long as $\Delta\mathcal{D}_{ijk}>0$ for any relevant $i$, $k$ and $j$. Notably, the set $j$ is taken from, $\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}$, is the set of possible hypotheses to participate in the Armitage Test against $H_{i}$’s representative. Finally, by averaging over the underlying hypothesis using the priors $\{\pi_{i}\}_{i=0}^{H-1}$, we obtain that
$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$
as desired. All that remains is bounding the number of stages; There is at least one stage and at most $H-1$ stages, so
$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}.$
A-D Proof of Corollary 1
Combining Lemma 2 and Lemma 3 with (1) yields
$\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}\leq\frac{\delta}{H^{2}}\times\frac{H}{\varepsilon}\log\frac{H}{\delta}+\delta=\frac{1}{\varepsilon}\times\frac{1}{\frac{H}{\delta}}\log\frac{H}{\delta}+\delta$
, which vanishes when $\delta\to 0$. When $H\to\infty$, only the first addend vanishes, and we are left with $\delta$.

### v4

@@
 The DM collects observations according to its determined actions and knows the distributions under each hypothesis.
-We propose a deterministic and adaptive multi-stage hypothesis-elimination algorithm where the DM selects an action, applies it repeatedly, and discards hypotheses in light of its obtained observations.
+We propose a deterministic and adaptive multi-stage hypothesis-elimination strategy where the DM selects an action, applies it repeatedly, and discards hypotheses in light of its obtained observations.
 The DM selects actions based on maximal separation expressed by the distance between the parameter vectors of each distribution under each hypothesis.
 Close distributions can be clustered, simplifying the search and significantly reducing the number of required observations.
-Our algorithm achieves vanishing Average Bayes Risk (ABR) as the error probability approaches zero, i.e., the algorithm is asymptotically optimal.
-Furthermore, we show that the ABR cannot exceed the error probability when the number of hypotheses grows.
-Simulations are carried out to evaluate the algorithm’s performance compared to another multi-stage hypothesis-elimination algorithm, where an improvement of 5 to 6 orders of magnitude in the mean number of observations required for success is observed.
+Our algorithms achieve vanishing Average Bayes Risk (ABR) as the error probability approaches zero, i.e., the algorithm is asymptotically optimal.
+Furthermore, we show that the ABR is bounded when the number of hypotheses grows.
+Simulations are carried out to evaluate the algorithm’s performance compared to another multi-stage hypothesis-elimination algorithm, where an improvement of several orders of magnitude in the mean number of observations required is observed.
 I Introduction
-HT is a statistical method used to evaluate competing hypotheses and determine which one best fits a sampled population based on sample data.
-HT has been in use for centuries, long before formal statistical methods were developed, and has been applied across highly diverse domains.
-For example, it has been used to identify the physical model that best fits a set of measurements, such as determining the Earth’s circumference or verifying a quantum mechanics hypothesis.
-HT is also an integral part of modern scientific practice, particularly in medical research where it’s used to evaluate the effectiveness of treatments and diagnose conditions based on clinical tests (e.g., testing for infectious diseases like COVID-19).
 In recent years, HT has gained additional significance with the recent vast interest in Machine Learning (ML) and Artificial Intelligence (AI).
@@
 In classic binary HT, this is often accomplished by computing the LLR or its non-logarithmic equivalent, comparing it to a predetermined threshold, and selecting the corresponding hypothesis based on this comparison.
-In many practical scenarios, the DM needs to collect and process data in real-time, and often can select which samples or data sources to examine next.
+In many practical scenarios, the DM needs to collect and process data in real-time and often can select which samples or data sources to examine next.
 For example, a physician can determine which subsequent diagnostic tests a patient should undergo to identify their condition, or a network administrator can choose which routers to monitor to detect and classify potential cyber-attacks.
@@
 In this context, the DM declares the tournament winner as the underlying hypothesis.
+Armitage test’s asymptotic optimality has been studied in [4], and its expected number of observations is further studied in [5].
 Chernoff [6] extended the Wald test setup to incorporate actions, leading to active binary SHT.
@@
 Chernoff proposed a stochastic policy (i.e., actions are drawn according to some non-degenerate distribution) for selecting actions sequentially, allowing the DM to “shape” observations through chosen actions to optimize accuracy and detection delay.
-Chernoff’s stochastic approach was later fine-tuned by Bessler in [5] and by Nitinawarat et al.
-in [6] to encompass active multihypothesis testing like considered in this work.
-Solutions based on Dynamic Programming and Linear Fractional Programming also exist, e.g., [7, 8, 9].
-Deep Learning and Reinforcement Learning have also been considered, e.g., [10, 11, 12].
-A prime example of a deterministic action-selecting policy is the DGF policy used for anomaly detection in [13] and its numerous extensions, e.g., [14, 15, 16].
-In the DGF policy, one observes that the accumulated LLRs corresponding to infected and clean distributions are random walks growing apart when enough observations are collected and when the gap between the largest and second-largest LLRs is sufficiently large, the DGF policy stops and declares the hypothesis corresponding to the highest accumulated LLR as true.
-Although the DGF policy outperforms Chernoff’s approach, the main drawback of the DGF policy is that it is tailored for anomaly detection.
-Namely, it assumes actions produce observations from only two possible distributions.
-In this paper, we take a different approach from the previously mentioned works by adopting the elimination approach rather than the typical search approach.
-Namely, instead of looking for enough evidence to conclude which hypothesis is true, the obtained observations are utilized to identify which hypothesis is “definitely not true.” To this end, the DM leverages the ability of some actions to separate some hypotheses better than others to swiftly and accurately discard inconsistent hypotheses.
-In particular, the DM selects its current action based on maximal separation of the remaining hypotheses, applies it repeatedly to collect enough observations to infer which hypotheses are incorrect using the Armitage Test, and discards them before repeating the procedure unless a single hypothesis remains.
-Still, it is possible for the maximal separation approach to require the DM to collect a large number of observations, all the more so when the obtained observations follow highly similar distributions under different hypotheses.
-Thus, we suggest clustering these distributions and require the DM only to identify which cluster is correct and discard all incorrect hypotheses accordingly.
-The closest work to our approach is the recently proposed algorithm by Gan, Jia, and Li in [17, Algorithm 2], which we refer to as the GJL algorithm.
-In GJL, the action capable of minimizing the number of surviving hypotheses is selected in each stage, and the (fixed) number of observations is based on separating the two closest, yet not identical, observation distributions.
-This philosophy of minimal survivors directly contrasts our approach of fast elimination incited by the maximal separation policy.
-Assuming that the clustering is done properly, we show that our algorithm is asymptotically optimal in terms of vanishing Average Bayes Risk (ABR) when the desired error probability approaches zero and is bounded when the number of hypotheses grows.
+Other stochastic approaches can be found in the literature, including active multihypothesis models like those considered in this work, e.g., [7, 8, 9, 10].
+Several deterministic policies have been explored in recent years.
+Novel examples include incorporating ML or Deep Learning with the Wald Test, e.g., [11, 12].
+Other learning approaches directly optimize the number of samples without invoking the Wald Test, e.g., [13, 14].
+A prime example of non-learning action-selecting policy is the DGF policy used for anomaly detection in [15, 16, 17, 18].
+In the DGF policy, the DM tracks the gap between the highest and second-highest accumulated LLRs, and when it is sufficiently large, it delivers its decision.
+Although the DGF outperforms Chernoff’s approach, the main drawback of the DGF policy is that it is tailored for anomaly detection; that is, actions produce observations from only two possible distributions, and extension to multiple distributions per action model is not straightforward.
+In this paper, we depart from the previously mentioned conventional approaches by employing an elimination strategy rather than a traditional search method.
+Instead of accumulating evidence to identify the correct hypothesis, we use observations to systematically eliminate hypotheses that are “almost surely” incorrect.
+The rationale for preferring elimination over the search-for-winner strategy stems from a key limitation in the latter approach: in the search strategy, hypotheses compete against each other simultaneously, with the DM’s actions and sampling decisions being guided by the underlying hypothesis and its closest hypotheses.
+While this approach may be effective with a few hypotheses, it becomes inefficient when handling multiple hypotheses because separating two close hypotheses requires an enormous number of samples.
+In contrast, the elimination strategy efficiently discards incorrect hypotheses by focusing on those that are most distinctly different from each other, typically requiring far fewer samples.
+The strategy proceeds sequentially, with each new action determined by the remaining hypotheses.
+Once only two hypotheses remain, the DM must determine which of these candidates is correct, with the key advantage that one of them is almost surely the true hypothesis, rather than having to isolate an unknown correct hypothesis from among many possibilities.
+To expedite the elimination process, we propose a clustering mechanism that discards multiple hypotheses at once rather than eliminating them sequentially.
+Specifically, for each possible action, the DM clusters the potential hypotheses based on their properties.
+It then selects the action that ensures sufficient separation between clusters by maximizing the minimum distance between them.
+The chosen action is used to draw samples and discard at least one entire cluster.
+Notably, cluster sizes vary—some may contain a single hypothesis, while others may include multiple hypotheses.
+Additionally, the number of clusters is not predetermined and can range from as few as two to as many as the total number of hypotheses.
+This process repeats: selecting an action that best separates the remaining hypotheses into clusters, using the action to collect samples, and discarding at least one cluster along with all its hypotheses.
+This continues until only a single hypothesis remains, which is then declared correct.
+We prove that, with proper clustering, our algorithm is asymptotically optimal in terms of vanishing Average Bayes Risk (ABR) as the desired error probability approaches zero while remaining bounded as the number of hypotheses increases.
+To highlight the benefits of the proposed “elimination” approach, we compare our algorithm with the recently introduced algorithm by Gan, Jia, and Li in [19, Algorithm 2], which we refer to as the GJL algorithm.
+Unlike our approach, GJL follows a “competition” strategy, where a single hypothesis “wins” over all others.
+However, because some hypotheses may be indistinguishable under a given action—that is, they yield the same distribution—GJL selects the action that eliminates the maximum number of hypotheses while minimizing the number of overlapping ones.
+After selecting an action, a fixed number of observations is collected, determined by the number required to distinguish the two closest hypotheses that do not yield identical distributions.
+This process continues until only one hypothesis remains.
 Finally, we numerically evaluate our algorithm and observe a significant performance improvement compared to GJL.
-Specifically, our algorithm improves the mean number of observations required until termination (and consequently, the ABR) by 5 to 6 orders of magnitude for any desired error probability.
+Specifically, our algorithm improves the mean number of observations required until termination (and consequently, the ABR) by several orders of magnitude for any desired error probability.
 II System Model
@@
 The $\ell_{2}$-norm between two vectors ${\underline{x}}$ and ${\underline{y}}$ is denoted as $\|{\underline{x}}-{\underline{y}}\|_{2}$.
-For the Kullback-Leibler Divergence (KLD) between two distributions, $f$ and $g$, we write $\mathcal{D}_{KL}(f\|g)$.
-Unless explicitly specified (e.g., $\ln$), all logarithms in this manuscript are in base two.
-Throughout this paper, we adopt the Bachmann–Landau big-O asymptotic notation, $\operatorname{\mathcal{O}}\left(\cdot\right)$, as defined in [18, Chapter 3].
+The Kullback-Leibler Divergence (KLD) between two distributions, $f$ and $g$, is denoted as $\mathcal{D}_{KL}(f\|g)$.
+Unless explicitly specified (e.g., $\ln$), all logarithms are in base 2.
 II-B Model
-The system model consists of a single, active, DM capable of obtaining observations according to the actions taken.
+The system model consists of a single, active, DM capable of obtaining samples according to the actions taken.
 The DM is tasked with correctly identifying the underlying hypothesis out of a finite set of hypotheses =$\mathcal{H}=\{0,1,\dots H-1\}$.
@@
 Let $\mathcal{A}\subset\mathbb{N}$ with $|\mathcal{A}|<\infty$ be the set of all actions available to the DM.
-We assume that the observations are independent and follow the same distribution when action $a_{n}$ is taken, with the distribution parameters depending on the underlying hypothesis.
+We assume that the samples are independent and follow the same distribution when action $a_{n}$ is taken, with the distribution parameters depending on the underlying true hypothesis.
 Namely, assuming $H_{i}$ is the underlying hypothesis and that action $a_{n}\in\mathcal{A}$ is taken at time step $n$, the DM obtains $X_{n}\sim f_{a_{n}}(\cdot\ ;{\underline{\theta}}_{i}(a_{n}))$, where ${\underline{\theta}}_{i}(a_{n})\in\mathbb{R}^{M_{a_{n}}}$ is the (vectorized) distribution parameter under $H_{i}$ and $f_{a_{n}}$ is its Probability Density Function (PDF).
@@
 The DM is tasked to identify the correct hypothesis (say $H_{i}$) out of $H$ possible hypotheses.
-By taking action $a_{n}$ at time step $n$, the DM obtains an observation $x_{n}$ whose PDF is $f_{a_{n}}(\cdot\ ;{\underline{\theta}}_{i}(a_{n}))$.
+By taking action $a_{n}$ at time step $n$, the DM obtains a sample $x_{n}\sim f_{a_{n}}(\cdot\ ;{\underline{\theta}}_{i}(a_{n}))$.
 The alphabet and size of $x_{n}$ and ${\underline{\theta}}_{i}(a_{n})$ may depend on the action $a_{n}$.
-Like [17], we make three additional assumptions:
+To simplify notation, we write $\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))$ instead of $\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{j}(a)))$.
+Similar to [19], we make three additional assumptions:
   1. (A1)
-    (Seperation) For any $a\in\mathcal{A}$ and $i$, $j\in\mathcal{H}$, $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}$ is either 0 or greater than some $\alpha>0$.
-Furthermore, there is no $a$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}=0$ for all $i$, $j\in\mathcal{H}$.
-Note that $\alpha$ can be very small.
+    (Separation) For any action $a\in\mathcal{A}$, for any $i$, $j\in\mathcal{H}$, $\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))$ is either 0 or greater than some known $\alpha>0$.
+Furthermore, there is no $a$ with $\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))=0$ for all $i$, $j\in\mathcal{H}$.
   2. (A2)
-    For any action $a$ there are $0<c_{1}\leq c_{2}$ such that any $i,j\in\mathcal{H}$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}>0$ also have $c_{1}\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{j}(a))\|f_{a}(\cdot;{\underline{\theta}}_{i}(a)))\leq\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}\leq c_{2}\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{j}(a))\|f_{a}(\cdot;{\underline{\theta}}_{i}(a)))\leq c_{2}\beta$ for some $\beta$.
+    For any action $a$ there are $0<c_{1}\leq c_{2}$ such that $c_{1}\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))\leq\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}\leq c_{2}\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))\leq c_{2}\beta$ for any $i,j\in\mathcal{H}$ and for known $\beta>0$.
   3. (A3)
-    (Validity) For all $i,j\in\mathcal{H}$ with $i\neq j$, there is some $a\in\mathcal{A}$ with $\|{\underline{\theta}}_{i}(a)-{\underline{\theta}}_{j}(a)\|_{2}^{2}>0$.
-The first assumption ensures that there are no meaningless actions and that distribution parameters are separated in $\mathbb{R}^{M_{a}}$.
-The second is technical and allows us to simplify hypothesis comparison by using the $\ell_{2}$-norm on the parameters rather than computationally prohibitive metrics on distributions (e.g., Total Variation Distance [19, Chapter 13.1]).
-The third assumption assures that multi-stage hypothesis-eliminating algorithms can stop.
+    (Validity) For all $i,j\in\mathcal{H}$ with $i\neq j$, there is some $a\in\mathcal{A}$ with $\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))>0$.
+The first assumption ensures that there are no meaningless actions and that distributions are separated under each action.
+The second is technical and allows us to simplify hypothesis comparison by using the $\ell_{2}$-norm on the distribution parameters rather than computationally prohibitive metrics (e.g., Total Variation Distance [20, Chapter 13.1]).
+The third assumption assures that some multi-stage hypothesis-eliminating algorithms can stop.
 Let $\Phi$ be the source selection process generating the action sequence $\{A_{n}\}_{n=1}^{\infty}$.
@@
 The decision rule is given by $\Psi\in\mathcal{H}$, i.e., $\Psi=i$ implies that $H_{i}$ is declared as true.
-Like many other works in the literature, e.g., [4, 6, 13, 20, 16], we assume the DM pays $\delta\in(0,1)$ to acquire a realization of $X_{n}$, and we are interested in the asymptotic behavior when $\delta\to 0$.
-Additionally, to emphasize the gains from the hypothesis clustering (which will be discussed in the following section), we are also interested in the regimes when $H$ is not too small, and, accordingly, we modify the common Bayesian approach.
-Namely, instead of the unit penalty accepted in other works, we assign a penalty of $H^{2}$ for a wrong declaration.
-Let $\Gamma\triangleq(\Phi,\Psi)$ be an admissible strategy for the SHT aiming at minimizing the ABR, $\delta\mathbb{E}\left[N\right]+H^{2}\times p_{e}$, where $N$ is the number of observations upon algorithm termination (i.e., detection delay), and $p_{e}$ is the error probability of the strategy $\Gamma$.
+Like many other works in the literature, e.g., [6, 8, 15, 21, 18], we focus on the Bayesian approach.
+Namely, let $\Gamma\triangleq(\Phi,\Psi)$ be an admissible strategy for the SHT aiming at minimizing the ABR, $\delta\mathbb{E}\left[N\right]+H^{2}\times p_{e}$, where $N$ is the number of samples upon algorithm termination (i.e., detection delay), $\delta$ is the sample cost, and $p_{e}$ is the error probability of $\Gamma$.
 Equivalently, it is a strategy minimizing
@@
 Accordingly, the desired error probability and the mean detection delay must be balanced.
-The normalized cost, $\frac{\delta}{H^{2}}$, plays a role in damping the mean detection delay; Specifically, it vanishes in $\delta$ only if the mean detection delay is sublinear in $\frac{1}{\delta}$ and subquadratic in $H$.
-While the motivation behind the former choice is clear, the latter choice damps the number of samples per stage to be sublinear in $H$ in cases when $H-1$ stages are used in a multi-stage hypothesis-eliminating algorithm.
-III Observation Acquisition Policy
-In this section, we formalize the suggested algorithm to acquire observations and compute the underlying hypothesis using them while minimizing the ABR.
-Like many works in literature, in order to identify which hypothesis is correct or not, the DM leverages the Test (LLRT) where the collected samples up to time step $n$ are used to compute the accumulated LLR which is compared to some fixed threshold $\gamma>0$ we will discuss later.
+The normalized cost, $\frac{\delta}{H^{2}}$, plays a role in damping the mean number of samples; Specifically, it vanishes in $\delta$ only if the mean number of samples is sublinear in $\frac{1}{\delta}$ and subquadratic in $H$.
+III Sample Acquisition Policy
+In this section, we formalize the suggested strategy for sample acquisition to compute the underlying true hypothesis.
+Our strategy consists of two flavors; without and with clustering, both of which enjoy vanishing ABR (Corollary 1), have vanishing error probability (Lemma 1) and more importantly, require a low mean number of samples (Theorem 1) as we show in the following section.
+Notably, the mean number of samples is lower than GJL’s if both select the same actions.
+Like many works in literature, in order to identify which hypothesis is correct or not, the DM leverages the (two-threshold) Armitage Test, where the collected samples up to time step $n$ are used to compute the accumulated LLR for any pair of hypotheses and each is compared to fixed thresholds $\gamma>0$ and $-\gamma$, to be discussed later.
 Namely, if the DM uses only action $a$, the pair-wise per-action accumulated LLR is:
 $L_{ij}(a,n)$ $\triangleq\sum_{t=1}^{n}\log\frac{f_{a}(x_{t};{\underline{\theta}}_{i}(a))}{f_{a}(x_{t};{\underline{\theta}}_{j}(a))}.$
-For simplicity, we will drop $a$ and $n$ and write $L_{ij}$.
+We will drop $a$ and $n$ and write $L_{ij}$ when they are clear from context.
 In our context, $H_{i}$ wins against $H_{j}$ if $L_{ij}\geq\gamma$, loses against $H_{j}$ if $L_{ij}<-\gamma$, and competes against $H_{j}$ otherwise.
-We will show in the following section that $2^{-\gamma}$ controls the per-stage error probability of the algorithm.
 Generally, identifying the underlying hypothesis has two distinct approaches.
-The first is the classic and straightforward approach common in literature where the DM collects observations and computes $L_{ij}$ until some $i^{*}$ wins against all other hypotheses (e.g., [2, 5]).
-While this approach is simple and fast, it may fail when the environment outputs observations with identical distributions under many different hypotheses, or in other words, the case we are interested in solving.
-The second approach, tailored for the above case, is hypothesis eliminating where the DM identifies which hypotheses are “definitely not true” and discards them until a single hypothesis remains and is declared as the underlying hypothesis.
+The first is the classic and straightforward approach common in literature where the DM collects samples and computes $L_{ij}$ until some $i^{*}$ wins against all other hypotheses (e.g., [2, 7]).
+While this approach is simple and fast, it may fail when the environment outputs samples with identical distributions under many different hypotheses, or in other words, the case we are interested in solving.
+The second approach, tailored for the above case, is hypothesis elimination, where the DM identifies which hypotheses are inconsistent with the collected samples and discards them until a single hypothesis remains, which is declared as true.
 Notably, algorithms of the first approach can be converted to the second by discarding the losing hypotheses and retrying using a different set of actions, as we do in this work.
 Once an action is selected, eliminating some hypothesis is straightforward by finding at least one hypothesis that wins against it.
-Thus, a multi-stage algorithm based on the LLRT is sufficient for one-by-one hypothesis elimination in our case since there is always some action capable of separating at least one hypothesis from the others due to assumption (A3).
-Naturally, one-by-one elimination is inefficient but can be improved by the action selection policy.
-Namely, at each stage, the DM can greedily pick the action capable of maximizing the number of losing hypotheses like in GJL [17, Algorithm 2].
-However, such a policy may result in a large number of samples and, accordingly, in large ABR.
-We adopt a different policy with both efficient hypothesis elimination (i.e., eliminating a set of hypotheses in each stage) and vanishing ABR.
-In each stage, we compute the action maximizing the separation, thus minimizing the number of samples required to eliminate hypotheses and, finally, minimizing the ABR.
-To enhance the hypothesis elimination process, we further suggest clustering hypotheses (per action) by proximity, as we will elaborate soon.
-Here, the DM only needs to identify the correct cluster rather than a set of hypotheses whose output distribution is identical and, accordingly, may use fewer samples.
+Thus, a multi-stage strategy based on the Armitage Test is sufficient for one-by-one hypothesis elimination in our case since there is always some action capable of separating at least one hypothesis from the others due to assumption (A3).
+Naturally, one-by-one elimination is inefficient, but it can be improved by the action selection policy.
+Namely, at each stage, the DM can greedily pick the action capable of minimizing the number of winning hypotheses, like in GJL [19, Algorithm 2].
+However, such a policy may result in a large number of samples, and, accordingly, large ABR.
+Thus, in order to both efficiently eliminate hypotheses and achieve faster vanishing ABR, we suggest taking actions that maximize the separation, therefore minimizing the number of samples required to eliminate hypotheses.
+To further enhance the elimination process, we suggest clustering hypotheses (per action) according to the sample distribution proximity.
+Then, the DM only needs to identify the correct cluster and, accordingly, may use fewer samples.
 III-A Preprocessing Step - Per-Action Hypothesis Clustering
@@
 The main purpose of the hypothesis clustering is to allow the DM to refine its “bad actions,” characterized by small KLD between some of their output distributions (e.g., either zero or very close to $\alpha$ from Assumption (A1)).
-Namely, to tell apart between two or more very close hypotheses, the DM would require, on average, a large number of observations.
-Therefore, since these hypotheses are very similar, the DM can significantly reduce the number of the collected observations by simply identifying the set of similar hypotheses (e.g., by identifying a single member of these hypotheses as we do in this work) and using a different action with better separation on the found set in the next stage.
-In other words, the DM can improve its performance in terms of mean detection delay (and, consequently, obtain lower ABR) by clustering hypotheses by their informational proximity defined by a proximity parameter $\varepsilon$ set by the user.
-Hypothesis clustering by their proximity (i.e., selecting $\varepsilon$) and selecting a representative from each cluster are not trivial.
+Namely, to distinguish between two or more very close hypotheses, the DM would require, on average, a large number of samples.
+Therefore, since these hypotheses are very similar, the DM can significantly reduce the number of the collected samples by simply identifying the set of similar hypotheses (e.g., by identifying a single member of these hypotheses as we do in this work) and using a different action with better separation on the found set in the next stage.
+In other words, the DM can improve its performance in terms of mean detection delay (and, consequently, obtain lower ABR) by clustering hypotheses by their informational proximity defined by a proximity parameter $\varepsilon_{a}$ set by the user.
+Clustering hypotheses by their proximity (i.e., selecting $\varepsilon_{a}$) and selecting a representative from each cluster is not trivial.
 Thus, we first discuss a key property of a good hypothesis clustering mechanism.
 Assume action $a$ is repeatedly used and that $H_{i}$ is the underlying hypothesis.
-Now, we observe the normalized accumulated LLR between $H_{k}$ in $H_{i}$’s cluster against $H_{j}$ that is not in $H_{i}$’s cluster converges to
-$\frac{1}{n}L_{kj}$ $\to\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{k}(a))}{f_{a}(X;{\underline{\theta}}_{j}(a))}\middle|H_{i}\right]$
-$=\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{j}(a)))$
-$\quad\quad-\mathcal{D}_{KL}(f_{a}(\cdot;{\underline{\theta}}_{i}(a))\|f_{a}(\cdot;{\underline{\theta}}_{k}(a)))\triangleq\Delta\mathcal{D}_{ijk}$ (2)
-in probability from the Asymptotic Equipartition Property (AEP) property [21, Theorem 11.8.1].
-Since $\frac{1}{n}\gamma\to 0$ for any fixed $\gamma$, we observe that $H_{i}$ can be declared as true if and only if the difference in (2) is positive.
-Therefore, any $H_{k}$ clustered with $H_{i}$ should have its KLD (from $H_{i}$ to $H_{k}$) strictly lower than $H_{i}$’s KLD from any $H_{j}$ not clustered with $H_{i}$.
+Now, we observe the normalized accumulated LLR between $H_{k}$ in $H_{i}$’s cluster against $H_{j}$ that is not in $H_{i}$’s cluster satisfies
+$\frac{1}{n}L_{kj}$ $\xrightarrow{n\to\infty}\mathbb{E}\left[\log\frac{f_{a}(X;{\underline{\theta}}_{k}(a))}{f_{a}(X;{\underline{\theta}}_{j}(a))}\middle|H_{i}\right]\triangleq\Delta\mathcal{D}_{ijk}(a)$
+$=\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))-\mathcal{D}_{KL}(H_{i}(a)\|H_{k}(a))$ (2)
+by the Asymptotic Equipartition Property (AEP) [22, Theorem 16.8.1], when the convergence is almost sure convergence.
+Like before, we will drop $a$ when it is clear from context.
+If the difference in (2) is positive, then $H_{k}$ would win against $H_{j}$ when a sufficiently large number of samples is collected when $H_{i}$ is true.
+Therefore, any $H_{k}$ clustered with $H_{i}$ should have $\mathcal{D}_{KL}(H_{i}(a)\|H_{k}(a))<\mathcal{D}_{KL}(H_{i}(a)\|H_{j}(a))$ for any $H_{j}$ not clustered with $H_{i}$.
 Since KLD is asymmetric, $H_{i}$ can be in the same cluster as $H_{k}$, but not vice-versa.
@@
 Let $\operatorname{param}\left(a\right)\triangleq\{{\underline{\theta}}_{i}(a):i\in\mathcal{H}\}$ be the set of distribution parameters under action $a$.
-Let $\mathcal{B}_{\varepsilon}({\underline{\nu}})\triangleq\{{\underline{\theta}}\in\operatorname{param}\left(a\right):\|{\underline{\theta}}-{\underline{\nu}}\|_{2}^{2}\leq\varepsilon\}$ be the $\varepsilon$-neighborhood of ${\underline{\nu}}$ composed of points from $\operatorname{param}\left(a\right)$.
-We say that ${\underline{y}}$ is reachable from ${\underline{x}}$ if there are ${\underline{x}}_{1},{\underline{x}}_{2},\dots,{\underline{x}}_{r-1}$ such that ${\underline{x}}={\underline{x}}_{1}$, ${\underline{x}}_{j}\in\mathcal{B}_{\varepsilon}({\underline{x}}_{j+1})$ and, additionally we have ${\underline{x}}_{r-1}\in\mathcal{B}_{\varepsilon}({\underline{y}})$.
-Similarly, ${\underline{x}}$ and ${\underline{y}}$ are connected if ${\underline{x}}$ is reachable from ${\underline{y}}$ and vice-versa.
-Thus, we define the cluster of $H_{i}$ as $\mathcal{C}_{i}(a,\varepsilon)\subseteq\operatorname{param}\left(a\right)$ to be a non-empty subset such that (i) for any ${\underline{\nu}}$, if ${\underline{\nu}}$ is reachable from ${\underline{\theta}}_{i}$ then ${\underline{\nu}}\in\mathcal{C}_{i}(a,\varepsilon)$, and (ii) any ${\underline{\nu}},{\underline{\theta}}\in\mathcal{C}_{i}(a,\varepsilon)$ are connected.
+Let $\mathcal{B}_{\varepsilon_{a}}({\underline{\nu}})\triangleq\{{\underline{\theta}}\in\operatorname{param}\left(a\right):\|{\underline{\theta}}-{\underline{\nu}}\|_{2}^{2}\leq\varepsilon_{a}\}$ be the $\varepsilon_{a}$-neighborhood of ${\underline{\nu}}$ composed of points from $\operatorname{param}\left(a\right)$.
+We say that ${\underline{y}}$ is reachable from ${\underline{x}}$ if there are ${\underline{x}}_{1},{\underline{x}}_{2},\dots,{\underline{x}}_{r-1}$ such that ${\underline{x}}={\underline{x}}_{1}$, ${\underline{x}}_{j}\in\mathcal{B}_{\varepsilon_{a}}({\underline{x}}_{j+1})$ (for $21\leq j\leq r-$ 2) and, additionally we have ${\underline{x}}_{r-1}\in\mathcal{B}_{\varepsilon_{a}}({\underline{y}})$.
+Thus, $H_{i}$’s cluster, $\mathcal{C}_{i}(a,\varepsilon_{a})\subseteq\operatorname{param}\left(a\right)$, is a non-empty subset such that for any ${\underline{\nu}}$, if ${\underline{\nu}}$ is reachable from ${\underline{\theta}}_{i}$ then ${\underline{\nu}}\in\mathcal{C}_{i}(a,\varepsilon_{a})$.
 Now, we define
-$\operatorname{equiv}\left(i,a\right)$$\triangleq\left\{k\in\mathcal{H}\ :\ {\underline{\theta}}_{k}\in\mathcal{C}_{i}(a,\varepsilon)\right\}$
+$\operatorname{equiv}\left(i,a\right)$$\triangleq\left\{k\in\mathcal{H}\ :\ {\underline{\theta}}_{k}\in\mathcal{C}_{i}(a,\varepsilon_{a})\right\}$
 $\operatorname{repr}\left(\mathcal{U},a\right)$$\triangleq\left\{\min\{\mathcal{U}\cap\operatorname{equiv}\left(i,a\right)\}:i\in\mathcal{U}\right\}$
 to be the set of equivalent hypotheses to $H_{i}$ and the set of representatives taken from $\mathcal{U}\subseteq\mathcal{H}$, respectively.
-Note that the choice of the smallest indices is arbitrary.
+$\operatorname{repr}\left(\mathcal{U},a\right)$ will contain the set of contestants (out of $\mathcal{U}$) in each stage of our strategy.
+Note that the choice of the smallest indices is arbitrary, and that Assumption (A3) still holds (for $\mathcal{U}$) if $|\operatorname{repr}\left(\mathcal{U},a\right)|>1$.
 If no clustering is used, we define $\operatorname{equiv}\left(i,a\right)\triangleq\{j\in\mathcal{H}:{\underline{\theta}}_{j}(a)={\underline{\theta}}_{i}(a)\}$ instead.
-Remark: We assume that even after the clustering, assumption (A3) holds.
-The use of the $\ell_{2}$-norm allows us to use rich density-based clustering over $\mathbb{R}^{M_{a}}$ (e.g., [22, Chapter 5]) to cluster hypotheses by proximity.
+The use of the $\ell_{2}$-norm allows us to use rich density-based clustering over $\mathbb{R}^{M_{a}}$ (e.g., [23, Chapter 5]) to cluster hypotheses by proximity.
 Notably, the famous Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [24] is tailored to our needs and is sufficient for clustering.
-DBSCAN groups together points with many shared neighbors (specified by the MinPts parameter) within their $\varepsilon$-neighborhood and labels all remaining points as noise or outliers.
-Furthermore, DBSCAN guarantees that clusters are $\varepsilon$ apart from each other.
+DBSCAN groups together points with many shared neighbors (specified by the MinPts parameter) within their $\varepsilon_{a}$-neighborhood and labels all remaining points as noise or outliers.
+Furthermore, DBSCAN guarantees that clusters are $\varepsilon_{a}$ apart from each other.
 III-B Multi-Staged Algorithm
@@
 In Line 1, we initialize the “alive” hypotheses, $H_{\mathrm{alive}}$, to be the set of all hypotheses.
-The algorithm thins this set in each stage until one or no hypothesis remains.
-If a single hypothesis remains at the end of the procedure, it is declared as true, and an error is raised otherwise.
+The algorithm thins this set in each stage until one or no hypotheses remain.
+If a single hypothesis remains at the end of the procedure, it is declared true.
 We term Lines 3-11 as stages.
-In each stage, an action $a$ is computed (Line 3).
-Specifically, each action is associated with a non-zero minimal distance between each two alive hypotheses, and the action with the maximal minimum distance is selected for the current round.
-Line 4 computes the representatives of each cluster out of the alive hypotheses.
-Now, we apply action $a$ repeatedly until a winner against all hypotheses, $i^{*}$, is found (Lines 6-9).
-Then, we update $H_{\mathrm{alive}}$ to be the alive hypotheses in $i^{*}$’s cluster (Lines 10-11).
+At the beginning of each stage, the action $a$ with the largest minimum distance between the alive hypotheses (i.e., best separation) is selected (Line 3).
+Line 4 computes the cluster representatives out of the alive hypotheses.
+This set contains the contestants for the current stage.
+Action $a$ is repeatedly applied until a winner against all hypotheses, $i^{*}$, is found (Lines 6-9).
+Then, $H_{\mathrm{alive}}$ is updated to be the alive hypotheses in $i^{*}$’s cluster (Lines 10-11).
 Algorithm 1 Multi-Stage LLR-Based SHT
@@
 In this section, we analyze Algorithm 1 in terms of error probability, mean detection delay, and the scaling of the ABR as functions of $\delta$ and $H$.
-Notably, all previous results regarding the Armitage Test, e.g., vanishing error probability or bounded mean detection delay, do not trivially follow from using the Armitage Test at each stage due to the clustering, and accordingly, all computations must be modified.
-Due to brevity, the proofs are attached in Appendix A.
-We start with the error probability.
-To bound the error probability of the algorithm, we first bound the per-stage error probability.
-Each stage of the algorithm is terminated when some hypothesis has all of its LLRs exceed $\log\frac{H}{\delta}$.
-At first glance, there are two possible error events; the first is that the stage winner (at some stage) is not the representative of the real hypothesis, and the second is the event when two different hypotheses win in the current round.
-We argue that the latter case never occurs in the following lemma:
+Notably, previous results regarding the Armitage Test, e.g., vanishing error probability, do not trivially follow when clustering hypotheses, and accordingly, some computations must be modified.
+When no clustering is used, however, the previous results naturally follow.
+We start with stating the main result:
+Theorem 1.
+Let $\tau_{r}$ be the number of samples used in stage $r$ when using action $a_{r}$, $H_{\mathrm{alive}}^{(r)}$ be the alive hypotheses in stage $r$, and $k$ be the representative from $H_{i}$’s cluster.
+Then $\mathbb{E}\left[\tau_{r}|H_{i}\right]=(1+o(1))\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ for any $i$ where the little-O is with respect to $\delta\to 0$.
+Furthermore, $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\max_{a\in\mathcal{A}}\left\{\frac{H}{\varepsilon_{a}}\log\frac{H}{\delta}\right\}$.
+Proof:
+The result on $\mathbb{E}\left[\tau_{r}|H_{i}\right]$ is a direct result of applying [4, Theorem 4.1] over the set of hypotheses participating in stage $r$’s Armitage Test.
+Notably, the theorem is applicable only if $\Delta\mathcal{D}_{ijk}>0$, i.e., for an appropriate choice of $\varepsilon_{a}$.
+The bounds on $\mathbb{E}\left[N\right]$ follow since the number of stages is at least one and at most $H-1$.
+∎
+We argue that the non-hypothesis clustering flavor of our algorithm improves over GJL if both use the same action sequence.
+Let $\mathcal{GJL}(a,\mathcal{U})=\{(l,m)\in\mathcal{U}^{2}:\mathcal{D}_{KL}(H_{l}(a)\|H_{m}(a))>0\}$ be the set of pairs from a hypotheses subset $\mathcal{U}\subseteq\mathcal{H}$ whose KLD under action $a$ is not zero.
+Let $\mathcal{D}_{\min}(a,\mathcal{U})=\mathop{\operatorname{argmin}}_{(l,m)\in\mathcal{GJL}(a,\mathcal{U})}\left\{\mathcal{D}_{KL}(H_{l}(a)\|H_{m}(a))\right\}$ be the smallest non-zero KLD of hypotheses $\mathcal{U}$ under action $a$.
+Regardless of which hypothesis is true, if GJL applies action $a$, it uses $\tau_{GJL}(a,\mathcal{U})=\log\frac{H}{\delta}/\mathcal{D}_{\min}(a,\mathcal{U})$ samples (i.e., average worst case).
+This quantity should be compared to the mean number of samples per stage in Theorem 1, which is lower than GJL’s from the maximality of $\tau_{GJL}(a,\mathcal{U})$.
+Now, we observe that, conditioned on success, if both algorithms use the same action sequences, then both algorithms use the same alive hypotheses at each stage.
+Consequently, our algorithm uses fewer samples on average than GJL.
+Now, we shift our focus to show asymptotic optimality. Accordingly, we bound the error probability of our strategy, given in the following lemma:
 Lemma 1.
-Let $i_{r}^{*}$ be the winner at stage $r$. Then, $i_{r}^{*}$ is unique.
-Proof:
-See Appendix A-A. ∎
-Now that we have established that there are no ties, we bound the former error probability with the following lemma:
-Lemma 2.
-For any action $a$, there exists some $\eta_{a}\in[0,1)$ such that $\Delta\mathcal{D}_{ijk}>\eta_{a}$, the current action still can distinguish between at least two hypotheses, and for any $\varepsilon\leq\eta_{a}$, the per-stage error probability cannot exceed $\frac{\delta}{H}$.
+For any action $a$, there exists some $\eta_{a}\in(0,1)$ such that (i) $\Delta\mathcal{D}_{ijk}>\eta_{a}$.
+(ii) The current action can still distinguish between at least two hypotheses.
+(iii) for any $\varepsilon_{a}\leq\eta_{a}$, the per-stage error probability cannot exceed $\frac{\delta}{H}$.
 Particularly, $p_{e}\leq\delta$.
 Proof:
-See Appendix A-B. ∎
-Now, we focus on the scaling laws of the mean detection delay, $\mathbb{E}\left[N\right]$, given in the following lemma:
-Lemma 3.
-Let $\tau_{r}$ be the number of observations used in stage $r$ when using action $a_{r}$, $H_{\mathrm{alive}}^{(r)}$ be the alive hypotheses in stage $r$, and $k$ be the representative from $H_{i}$’s cluster.
-Then $\mathbb{E}\left[\tau_{r}|H_{i}\right]\approx\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ for any $i$.
-Furthermore, $\mathbb{E}\left[\tau_{r}|H_{i}\right]$ is bounded, i.e., $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}|H_{i}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$ and particularly, $\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}$.
+The detailed proof can be found in Appendix A-A.
+We first bound the probability for $H_{j}$ to win against $H_{i}$ when $H_{i}$ true but is represented by $H_{k}$ by $\frac{\delta}{H}\times\kappa_{aijk}$ for some $\kappa_{aijk}$ utilizing a similar computation made by Chernoff in [6].
+Then, we leverage the clustering to show that there exists some $\eta_{a}$ such that $\kappa_{aijk}\leq 1-\eta_{a}$.
+Any $\varepsilon_{a}\leq\eta_{a}$ ensures the same, so the claim on the per-stage error probability follows.
+Once the per-stage error probability bound is established, the result on $p_{e}$ immediately follows from the union bound.
+∎
+Corollary 1 (Asymptotic Optimality).
+For a suitable choice of $\{\varepsilon_{a}\}_{a\in\mathcal{A}}$, Algorithm 1 has vanishing ABR when $\delta\to 0$. Furthermore, the ABR does not exceed $\delta$ when $H\to\infty$.
 Proof:
-See Appendix A-C. ∎
-Remark: Although $\mathbb{E}\left[N\right]$ seems identical to non-clustering cases, it is, in fact, smaller as the expectation not only contains fewer positive addends, but each addend is also smaller.
-Corollary 1 (Asymptotic Optimality).
-For a suitable choice of $\varepsilon$, Algorithm 1 has vanishing ABR when $\delta\to 0$. Furthermore, the ABR does not exceed $\delta$ when $H\to\infty$.
-Proof:
-See Appendix A-D. ∎
+Let $\varepsilon\triangleq\min_{a\in\mathcal{A}}\{\varepsilon_{a}\}$.
+Combining Lemma 1 and Theorem 1 yields $\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}\leq\frac{\delta}{H^{2}}\times\frac{H}{\varepsilon}\log\frac{H}{\delta}+\delta=\frac{1}{\varepsilon}\times\frac{1}{\frac{H}{\delta}}\log\frac{H}{\delta}+\delta$ , which vanishes when $\delta\to 0$.
+When $H\to\infty$, only the first addend vanishes, and we are left with $\delta$.
+∎
 V Numerical Results
-In this section, we present numerical results illustrating the performance of our algorithm compared to GJL.
+In this section, we present numerical results illustrating the performance of our algorithms (with and without clustering) compared to GJL.
 To this end, we have conducted simulations.
-Namely, we compare GJL to our algorithm in two scenarios: (i) when no clustering is used.
-(ii) clustering is used with a suitable proximity parameter.
 In the simulations, we have used uniform priors, i.e., $\pi_{i}=\frac{1}{H}$.
@@
 All random means were drawn only once.
-For the clustering algorithm, DBSCAN, we use $\text{minPts}=1$ and proximity parameters $\varepsilon\in\{0,0.1\}$.
-In Figure 2, we present the results for the two cases.
-The dashed lines present the ABR for the case where all actions produce observations following the normal distribution, whereas non-dashed lines present the ABR for the case where all actions produce exponentially distributed observations.
+For the clustering algorithm, DBSCAN, we use $\text{minPts}=1$ and (manually tuned) proximity parameter $\varepsilon_{a}=\varepsilon=0.1$ for any $a$.
+The simulation results are presented in Figure 2.
+The dashed lines present the ABR for the case where all actions produce normally distributed samples, whereas the non-dashed lines present the ABR for the case where all actions produce exponentially distributed samples.
 The orange curves correspond to GJL, and the blue and green curves correspond to our algorithm without and with clustering, respectively.
-Each algorithm attains vanishing ABR as $\delta\to 0$, but the non-greedy selections made by each instance of our algorithm enjoys significantly lower ABR, which is attributed to lower mean detection delay.
+Each algorithm attains vanishing ABR as $\delta\to 0$, but the non-greedy selections made by each instance of our algorithm enjoy significantly lower ABR, attributed to a lower mean number of samples.
 […]
 Figure 2: Comparison of the ABR (equation (1)) for $H=16$ using our algorithm (with proximity parameter $\varepsilon\in\{0,0.1\}$) against GJL.
-The observations obtained are either normally distributed (dashed) or exponentially distributed (non-dashed).
+The samples obtained are either normally distributed (dashed) or exponentially distributed (non-dashed).
 The ABR vanishes as $\delta\to 0$ for all instances, but GJL’s greedy selection has a slower decay.
@@
 A-A Proof of Lemma 1
+At first glance, there are two possible error events; the first is that the stage winner (at some stage) is not the representative of the true hypothesis, and the second is the event when two different hypotheses win in the current round.
+We argue that the latter case never occurs in the following proposition:
+Proposition 1.
+Let $i_{r}^{*}$ be the winner at stage $r$. Then, $i_{r}^{*}$ is unique.
+Proof:
 Assume to the contradiction that there are two winners, $i_{r}^{*}$ and $j_{r}^{*}$.
@@
 However, due to symmetry, we have $\log\frac{H}{\delta}\leq L_{i_{r}^{*}j_{r}^{*}}=-L_{j_{r}^{*}i_{r}^{*}}\leq-\log\frac{H}{\delta}$, contradiction.
-A-B Proof of Lemma 2
+∎
+Now that we have established that there are no ties, we are ready to bound the former error probability.
 Since the number of alive hypotheses can only decrease in each stage, there can be no more than $H$ stages.
@@
 Consequently, $\frac{\delta}{H}$ bounds the per-stage error probability, and as the section preamble suggests, the error probability of the algorithm cannot exceed $\frac{\delta}{H}\times H=\delta$.
-A-C Proof of Lemma 3
-Much like in the proof of the previous lemma, we notice that the Armitage Test essentially conducts $\operatorname{\mathcal{O}}\left(H^{2}\right)$ Wald Tests in parallel using the same observation vector.
-Due to the algorithm’s stage termination condition when clustering, we are interested in $\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ and the number of observations required for it to exceed $\log\frac{H}{\delta}$, where $k$ is the index of the hypothesis representing $H_{i}$ like before.
-Using Wald’s Identity, we have
-$\mathbb{E}\left[L_{kj}|H_{i}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{\tau_{r}}\log\frac{f_{a_{r}}(x_{t};{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(x_{t};{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
-$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\mathbb{E}\left[\log\frac{f_{a_{r}}(X;{\underline{\theta}}_{k}(a_{r}))}{f_{a_{r}}(X;{\underline{\theta}}_{j}(a_{r}))}\middle|H_{i}\right]$
-$=\mathbb{E}\left[\tau_{r}|H_{i}\right]\times\Delta\mathcal{D}_{ijk}$ (5)
-for any relevant $i$, $k$, and $j$ (and particularly, for the LLR minimizer). Denote $\hat{p}_{e}$ as the per-stage error probability. From the stopping rule, we have
-$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]=\hat{p}_{e}\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\exists j\neq i\text{ wins}\right]$
-$+(1-\hat{p}_{e})\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i},\min_{l}L_{kl}\geq\log\frac{H}{\delta}\right]$
-using Bayes.
-Now, in similar arguments made by Wald in [1, Chapter 4], we approximate $\hat{p}_{e}\approx\frac{\delta}{H}$ for small enough $\delta$.
-Furthermore, conditioned on success, we approximate $\min_{l}L_{kl}\approx\log\frac{H}{\delta}$ for small enough $\delta$.
-On the other hand, conditioned on error, we observe that $\min_{l}L_{kl}<\log\frac{H}{\delta}$ since otherwise there would be no error as the winner is unique per Lemma 1.
-This inequality can be weakened, and we obtain that
-$\mathbb{E}\left[\min_{l}L_{kl}\middle|H_{i}\right]$ $\lesssim\left(1-\frac{\delta}{H}\right)\log\frac{H}{\delta}+\frac{\delta}{H}\log\frac{H}{\delta}$
-$=\log\frac{H}{\delta}.$ (6)
-Specifically, combining (5) and (6) yields
-$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\lesssim\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
-$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$ (7)
-and the upper bound follows.
-Finally, since $i$ and $j$ are not clustered together, $\Delta\mathcal{D}_{ijk}\geq\varepsilon$ and we bound (7) by $\frac{1}{\varepsilon}\log\frac{H}{\delta}$.
-The lower bound
-$\mathbb{E}\left[\tau_{r}|H_{i}\right]$ $\geq\frac{\log\frac{H}{\delta}}{\min_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\Delta\mathcal{D}_{ijk}}$
-$=\max_{j\in\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}}\left\{\frac{\log\frac{H}{\delta}}{\Delta\mathcal{D}_{ijk}}\right\}$
-$\geq\frac{\log\frac{H}{\delta}}{\beta}$
-follows from applying [24, Theorem 3.1] (using the law of large numbers as in (2)) as long as $\Delta\mathcal{D}_{ijk}>0$ for any relevant $i$, $k$ and $j$.
-Notably, the set $j$ is taken from, $\operatorname{repr}\left(H_{\mathrm{alive}}^{(r)},a_{r}\right)\setminus\{k\}$, is the set of possible hypotheses to participate in the Armitage Test against $H_{i}$’s representative.
-Finally, by averaging over the underlying hypothesis using the priors $\{\pi_{i}\}_{i=0}^{H-1}$, we obtain that
-$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[\tau_{r}\right]\leq\frac{1}{\varepsilon}\log\frac{H}{\delta}$
-as desired. All that remains is bounding the number of stages; There is at least one stage and at most $H-1$ stages, so
-$\frac{1}{\beta}\log\frac{H}{\delta}\leq\mathbb{E}\left[N\right]\leq\frac{H}{\varepsilon}\log\frac{H}{\delta}.$
-A-D Proof of Corollary 1
-Combining Lemma 2 and Lemma 3 with (1) yields
-$\frac{\delta}{H^{2}}\mathbb{E}\left[N\right]+p_{e}\leq\frac{\delta}{H^{2}}\times\frac{H}{\varepsilon}\log\frac{H}{\delta}+\delta=\frac{1}{\varepsilon}\times\frac{1}{\frac{H}{\delta}}\log\frac{H}{\delta}+\delta$
-, which vanishes when $\delta\to 0$. When $H\to\infty$, only the first addend vanishes, and we are left with $\delta$.